Please solve this qs for me show all steps and clear explaining

At point A the wooden block below is traveling at
Answer #1


a)

work done by frictionWf = fk*AB*cos180

fk = uk*m*g


AB = 0.8 m

Wf = -0.22*1.8*9.8*0.8 = -3.10464 J

(b)

work done by gravity Wg = m*g*h = 1.8*9.8*1.4 = 24.696 J

(c)

work energy theorem


total work done = change in KE


W = 0.5*m*(v^2-u^2)

24.696 - 3.10464 = 0.5*1.4*(v^2-2.3^2)


v = 6.01 m/s   <<<<<-----------answer